40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Sep 13 09:15:15 2021, Paco, Update, LSC, MC unlocked? VAC_2021-09-13_09-32-45.png
    Reply  Mon Sep 13 15:14:36 2021, Anchal, Update, LSC, Xend Green laser injection mirrors M1 and M2 not responsive 
       Reply  Thu Sep 30 14:13:18 2021, Anchal, Update, LSC, HV supply to Xend Green laser injection mirrors M1 and M2 PZT restored 
Message ID: 16322     Entry time: Mon Sep 13 15:14:36 2021     In reply to: 16320     Reply to this: 16368
Author: Anchal 
Type: Update 
Category: LSC 
Subject: Xend Green laser injection mirrors M1 and M2 not responsive 

I was showing some green laser locking to Tega, I noticed that changing the PZT sliders of M1/M2 angular position on Xend had no effect on locked TEM01 or TEM00 mode. This is odd as changing these sliders should increase or decrease the mode-matching of these modes. I suspect that the controls are not working correctly and the PZTs are either not powered up or not connected. We'll investigate this in near future as per priority.

ELOG V3.1.3-